I'm having trouble finding the answer
solve: 2x^2 + 2x +10=0

Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

2x^2 + 2x +10 = 0
:
divide both sides of = by 2
:
x^2 +x +5 = 0
:
solve by completing the square
:
x^2 +x + 1/4 = -5 + 1/4
:
(x + 1/2)^2 = -19/4
:
take square root of both sides of =
:
x + 1/2 = square root(-19/4)
:
the solution is two complex numbers
:
*************************************
x = -i/2 * (square root(19) - i)
x = i/2 * (square root(19) + i)
:
note i = square root(-1) and i^2 = -1
:
**************************************
:

Answer by Alan3354(69443)

solve: 2x^2 + 2x +10=0
----------
I would use the quadratic equation.
---
Completing the square works, too.